OpenStudy (anonymous):

Who wants a medal? State the domain and range of y=-3/4x+4!!

OpenStudy (anonymous):

\[y=-3/4x+4\]

OpenStudy (zarkon):

for \[\frac{-3}{4x+4}\] the domain is all reals except x=-1

OpenStudy (anonymous):

oh no it's not for that. i'm not sure how to type it in though! I'm new here -____-

OpenStudy (zarkon):

\[\frac{-3}{4x}+4\]

OpenStudy (anonymous):

it's y= -3 over 4x +4. it's in the y=mx+b form.

OpenStudy (anonymous):

i mean -3 over 4, then x.

OpenStudy (zarkon):

ah \[\frac{-3}{4}x+4\]

OpenStudy (anonymous):

yes! (:

OpenStudy (zarkon):

all reals

OpenStudy (anonymous):

thank you!